OpenStudy (anonymous):

What is the graph of the inequality? –4y ≥ 8x – 2 (1 point)

OpenStudy (anonymous):

yikes

OpenStudy (anonymous):

\[4y\leq -8x+2\] \[y\leq -2x+\frac{1}{2}\]

OpenStudy (anonymous):

not b or c right

OpenStudy (anonymous):

slope is negative, shade below

OpenStudy (anonymous):

so actually it would be B or C because those are the ones with the negative slope pick B because it is shaded below

OpenStudy (anonymous):

oops hehe. ok i think b

OpenStudy (anonymous):

good done?

OpenStudy (anonymous):

4 more questions